OpenStudy (anonymous):

Help understanding how to find the answer please. (a + 3)(a - 2)

OpenStudy (anonymous):

Know FOIL method?

OpenStudy (anonymous):

I don't

OpenStudy (anonymous):

F= first O= over I = inner L= last So multiply the first values with each other (a x a) What does this give?

OpenStudy (anonymous):

After this multiply the same a with the (-2) now in the second parenthesis. What does this give? (a x -2) = ?

OpenStudy (anonymous):

Above was the " Over" part.

OpenStudy (anonymous):

ok im lost

OpenStudy (anonymous):

what does a x a give?

OpenStudy (anonymous):

a2?

OpenStudy (anonymous):

a^2 right?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

yes. Write this down

OpenStudy (anonymous):

and 3*2=6

OpenStudy (anonymous):

look at your equation as we multiply. We just multiplied your first a with the second a.

OpenStudy (anonymous):

You skip 2 multiplication steps darling... Lets do them together.

OpenStudy (anonymous):

oh oops, ok

OpenStudy (anonymous):

Remember First, over, inner, last! (FOIL) Thus the "first" is a x a = a^2 The " over" is a x -2 = -2a The " inner" is 3 x a = 3a The " Last" is 3 x -2 = -6

OpenStudy (anonymous):

Thus you get: a^2 - 2a + 3a -6. (What can you do now to simplify this equation further?)

OpenStudy (anonymous):

so what I need to do is when I get to the part including the number, multiply each number by the letter??

OpenStudy (anonymous):

My fill in the blank for this equation is this (___-a_____)

OpenStudy (anonymous):

Take the first part : (a + 3) now just work with a first ( forget the +3 for now) and multiply it with your second parenthesis: (a -2) Thus a x a and a x -2

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

Anyways the answer is a^2 + a - 6. Read my instructions carefully and try to do it yourself again.

OpenStudy (anonymous):

would I do the same for this on : (x^2 + 4)(x^2 - 4)

OpenStudy (anonymous):